Fill in the blanks in the table below. Assume that MPC is constant for everyone in the economy.

MPC

Spending Multiplier

Change in government spending

Change in income

100

$15

20

$2,000

0.6

-$400

0.5

$450

$900
Answer #1

Answer

MPC=1-MPS
MPS=1/spending multiplier

Working as per colum

MPS=1/100=0.01
MPC=1-.01=0.99

=====
MPS=1/20=0.05
MPC=1-0.05=0.95


=====
Spending multiplier=1/(1-MPC)
MPC=0.6
Spending multiplier =1/(1-0.6)=2.5
MPC=0.5
Spending multiplier =1/(1-0.5)=2
=======
Change in govenment spending =change in income /spending multiplier

Change in govenment spending =2000/20=$10
=========
Change in income =change in govenment spending *multiplier
change innincome for first balnk =15*100=1500
change inn income for second blank =-400*2.5=-$1000